Hi,
I know the answer 12 correct and 20 wrong, but I need a formular so that I can teach my son how I got the answer. Can you please give me a formula
To encourage John to work harder in math his mother said she would pay him 10 cents for each right answer and subtract 5 cents for each wrong answer. If he earned
20 cents after doing 32 problems, how many problems did John get right? How many did he get wrong? How many would he have to get right to earn more than a dollar?
Cheers,
Sureshni

Answer by Boreal(15235)

number of right answers=x
number of wrong answers are 32-x
10x-5(32-x)=20
10x-160+5x=20
15x=180
x=12; he made 120
32-x=20; he lost 100. Total is 20 cents.
----------------------------
to earn more than a dollar
10x-5(32-x)>100
10x-160+5x>100
15x > 260
x > 17 (18 will work but 17 won't)
17 should not quite work
17*10=170
15*5=85
The result is 85
18 will be 180-70=110